CARO Departmental stores sends sales reps to call on three kind of customers: Retail, Industrial and Professional. Sales revenue of $8,000 results from calling retail customer, $14,000 from calling Industrial customer and $20,000 from calling Professional customer. Each quarter a total of 3200hrs of sales reps time is available for calling on customer and $10,000 available on travelling expenses. Six hours of travel and selling are required to call on retail customer with $10 sales rep traveling expenses; 11hours for Industrial customer with $14 sales rep traveling expenses; and 25 hours for Professional customer with $25 sales rep traveling expenses. What is optimal customer mix for each quarter?
1
Expert's answer
2020-11-06T10:09:47-0500

Let number of retail clients be "r" , indurstrial

clients "i" and professional "p"


Constraints are time "t \le 3200" and cost "c \le 10,000"


The total time is shared as follows


"6r+11i+25p \le 3200"

The total cost is shared as follows


"10r+14i+25p \le 10,000"


Revenues are:


"8,000r+14,000i+20,000p"

Profits are the sum of revenue minus costs


"(8,000r-10r) + (14,000i-14i)+(20,000p-25p)"

"=(7,990r) + (13,986i)+(19,975p)"


The goal is to get the right mix of customers to maximize profits


Using Excel solver:





The optimal mix of customers that maximizes profits is 526 retail clients and 4 indurstrial clients.